OpenStudy (gavin39):

Can someone please answer this for me. I give medals. How can one fourthx − 3 = one halfx + 8 be set up as a system of equations? 4y + 4x = −12 2y + 2x = 16 4y − x = −12 2y − x = 16 4y + x = −12 2y + x = 16 4y − 4x = −12 2y − 2x = 16

OpenStudy (gavin39):

really need an answer, if anyone could please tell me.

OpenStudy (anonymous):

Strange question. You have only a single variable, and you have two functions that equal one another. The system required to solve this is exactly what you have. \[\frac{1}{4}x - 3 = \frac{1}{2}x+8\] I suppose you could set each equal to "y" and form them that way: \[y = \frac{1}{4}x - 3\] \[y= \frac{1}{2}x+8\] Which you can then rearrange to: \[y - \frac{1}{4}x = -3 \rightarrow 4y-x=-12\] \[y-\frac{1}{2}x = 8 \rightarrow 2y-x = 16\]